Listed below are the genome sizes (along with the total number of genes) for various organisms. Which genome has the largest gene density?

a. 68 Mb genome with 15,000 genes

b. 60,000 kb genome with 12,000 genes

c. 4.1 Mb genome with 3,500 genes

d. 20 Mb genome with 1,200 genes

e. 200 Mb genome with 14,300 genes
Answer #1

Answer: (c) 4.1 Mb genome with 3,500 genes

Explanation: The gene density of a genome is the ratio of the number of genes per number of base pairs, usually represented in million base pairs or megabase (Mb).

Calculating gene density for each set of given genome data:

Option (a) 68 Mb genome with 15,000 genes

Gene Density = 15000 / 68 = 220.58 genes/Mb

Option (b) 60,000 kb or 60 Mb genome with 12,000 genes

Gene Density = 12000 / 60 = 200 genes/Mb

Option (c) 4.1 Mb genome with 3,500 genes

Gene Density = 3,500 / 4.1 = 853.65 genes/Mb

Option (d) 20 Mb genome with 1,200 genes

Gene Density = 1,200 / 20 = 60 genes/Mb

Option (e) 200 Mb genome with 14,300 genes

Gene Density = 14,300 / 200 = 71.5 genes/Mb

Hence, Option (c) a genome with 853.65 genes/Mb has the largest gene density.

.

If you like my answer, please do not forget to thumbs up. I would be happy to discuss your concerns in the comment section below. Thank you!